a saturate solution of slightly soluble salt MA3 is 1.4 x 10^-3 M in MA3, what is Ksp value?
Answer #1

MA3 <----> A+3 + 3B-

S                 S        3S

Ksp = S*(3S)^3

       = 27S^4

       = 27*(1.4*10^-3)^4

       = 1.04*10^-10